OpenStudy (anonymous):

DESPERATE GEOMETRY HELP Determine if triangle DEF with coordinates D (2,1), E (3,5), and F (6,2) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changed could be made to make it equilateral?

OpenStudy (anonymous):

guess we gotta use the distance formula a few times to see if the distances are the same

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

as you can see it is an isosceles acute triangle not equilateral

OpenStudy (anonymous):

the distance between \((2,1)\) and \((3,5)\) is \[\sqrt{(3-2)^2+(5-1)^2}=\sqrt{1+16}=\sqrt{17}\]

OpenStudy (anonymous):

the distance between \((2,1)\) and \((6,2)\) is also \(\sqrt{17}\) the computation is almost identical

OpenStudy (anonymous):

but the distance between \((3,5)\) and \((6,2)\) is \[\sqrt{(6-3)^26(5-2)^2}=\sqrt{3^2+3^2}=3\sqrt2\]

OpenStudy (anonymous):

wouldnt you do 3 squared nd 3 squared and add thos and THEN square root it?

OpenStudy (anonymous):

hello??

OpenStudy (anonymous):

yes lets check and make sure it is clear you get the same thing i wrote since i guess i skipped a step

OpenStudy (anonymous):

\[\sqrt{3^2+3^2}=\sqrt{9+98}=\sqrt{18}=\sqrt{9\times 2}=3\sqrt2\]

OpenStudy (anonymous):

damn typo second one should be \(\sqrt{9+9}\)

OpenStudy (anonymous):

okiiee

OpenStudy (anonymous):

so how would you make it equilateral?

OpenStudy (anonymous):

hmm lets go back to the picture i guess we have to adjust one of the points

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

lol actually i have no idea let me think about it for a moment, it is kind of a weird question

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lol i can't think of what to do repost the second part, maybe someone has a good idea